A) 7a+3z=12
B) 3a+7z=6
Given the formulas above, what is the value of a-z?

Answer by nvegupta(2)

7a+3z=12
7a=12-3z
a=(12-3z)/7
3a+7z=6
3a=6-7z
a=(6-7z)/3
a=a
(12-3z)/7=(6-7z)/3
3(12-3z)=7(6-7z)
36-9z=42-49z
-9z+49z=42-36
40z=6
z=6/40
z=3/20
PUT VALUE OF z IN ANY SUITABLE EQUATION TO FIND VALUE OF a.
7a+3z=12
7a+3*3/20=12
7a+9/20=12
7a=12-9/20
7a=231/20
a=(231/20)/7
a=231/140
a-z
=(231/140)-(3/20)
=210/140
=21/14
=3/2